OpenStudy (anonymous):

Solve y' + 2xy = 2x^3, y(0) = 1.

OpenStudy (experimentx):

\[ \huge y = \frac{\int2x^3(e^{\int2x dx})dx +c}{e^{\int2x}}\]

myininaya (myininaya):

\[y'+2xy=2x^3 \] Multiply both sides by what is called the integrating factor It is already in this form: \[y'+p(x)y=q(x)\] The integrating factor is : \[e^{ \int\limits p(x) dx}\] So for this problem we have the integrating factor is: \[e^{\int\limits p(x) dx} =e^{\int\limits 2x dx}=e^{x^2}\] So the reason we multiply by the integrating factor is because it allows us to write the right hand side as \[(ye^{x^2})'\] Since this equals : \[y'e^{x^2} +2xye^{x^2} \text{ By product rule }\] Which is what we will have after multiplying both sides by the integrating factor \[e^{x^2}y'+e^{x^2}2xy=e^{x^2} 2x^3\] So we have again by product rule we can rewrite the right hand " \[(ye^{x^2})'=e^{x^2} 2x^3\] Now integrate both sides :) Don't forget +C after integrating